# Suiren-1.0 Technical Report: A Family of Molecular Foundation Models

### 23 Mar 2026 v1

@@ 5.1.1 MoleHB
We introduce Molelecular handbook (MoleHB), a comprehensive molecular property prediction benchmark encompassing 40+ heterogeneous tasks. The benchmark spans several critical scientific domains, including safety, structural, critical and saturation, energetic, thermal, solution, transport and fluctuation properties. All data points are sourced from [yaws1999chemical] and have been rigorously validated via wet-lab experiments to ensure high-fidelity, stable values. We propose two evaluation protocols: (1) Random split: a standard random split to evaluate performance under similar data distributions; and (2) Scaffold split: a strategy where molecules with larger atom counts are assigned to the validation set to assess the model’s structural extrapolation capabilities. Both the datasets and splitting protocols have been open-sourced to facilitate reproducible research. Results of Scaffold split are shown in Appendix C.
@@ 5.2.1 MoleHB (Random Split)
critical temperature K 27.3807/0.8340 9.3277/0.9787 30.6705/0.8850 7.2302/0.9775 22.49
critical pressure bar 2.0733/0.9045 0.7332/0.9765 2.9769/0.9084 0.5661/0.9792 22.79
critical density g/ml 0.0134/0.9253 0.0054/0.9814 0.0115/0.9497 0.0043/0.9786 20.37
critical volume cm3/mol 61.8034/0.8484 8.6329/0.9981 7.0201/0.9989 4.2825/0.9992 39.00
critical compressibility cm3 0.0189/0.6178 0.0066/0.8858 0.0098/0.8808 0.0057/0.8976 13.64
@@
flash point F 19.6978/0.8017 9.9274/0.9472 13.9535/0.9438 9.1004/0.9564 8.33
lower explosive limit vol% 0.1689/0.6854 0.0974/0.8926 0.1106/0.8341 0.0963/0.8561 1.13
upper explosive limit vol% 1.4188/0.5846 1.0460/0.6979 1.0023/0.6942 0.8346/0.7801 16.73
@@
heat capacity of liquid J/mol/K 16.0687/0.7922 5.9321/0.9812 5.8939/0.9792 4.5252/0.9815 23.22
heat capacity of solid J/mol/K 49.2807/0.9499 6.5306/0.9988 3.2724/0.9990 1.1395/0.9997 65.18
coefficient of thermal expansion of liquid 1/K 0.0002/0.1716 0.0001/0.8158 0.0001/0.6211 0.0000/0.9178 100.00
heat capacity of gas J/mol/K 24.1482/0.8849 4.4112/0.9685 3.2478/0.9691 2.7546/0.9690 15.19
@@
octanol water partition coefficient 0.3786/0.9414 0.1491/0.9877 0.1386/0.9858 0.1353/0.9878 2.38
solubility in water ppm(wt) 0.2110/0.9422 0.0564/0.9973 0.0456/0.9980 0.0370/0.9984 18.86
solubility in water containing salt ppm(wt) 0.2535/0.7540 0.0615/0.9924 0.0406/0.9961 0.0374/0.9967 7.88
solubility parameter (J/cm3)1/2 0.7111/0.8331 0.3873/0.9236 0.7614/0.8282 0.3600/0.9284 7.05
henrys law constant for compound in water 0.2856/0.9581 0.1517/0.9816 0.1930/0.9744 0.1298/0.9841 14.44
henrys law constant for gas in water 0.5454/0.3000 0.3017/0.7939 0.1773/0.9430 0.5104/0.4226 -187.87
@@
density of liquid g/ml 0.0331/0.9544 0.0160/0.9761 0.0145/0.9763 0.0120/0.9691 17.24
melting point K 19.9979/0.8199 14.1449/0.9053 24.7538/0.8050 11.5566/0.9245 18.30
boiling point K 19.8744/0.8891 6.2177/0.9906 12.3314/0.9777 4.5883/0.9928 26.21
vapor pressure mmHg 0.4837/0.5674 0.1596/0.9579 0.2000/0.9370 0.1505/0.9473 5.70
@@
liquid volume cm3/mol 19.5900/0.8583 3.9679/0.9945 8.9388/0.9876 2.1627/0.9972 45.50
acentric factor omega 0.0702/0.7591 0.0233/0.9293 0.0225/0.9312 0.0178/0.9019 20.89
refractive index - 0.0277/0.4260 0.0057/0.9314 0.0079/0.9276 0.0050/0.9351 12.28
radius of gyration $10^{-10}$m 0.3573/0.6317 0.1230/0.9715 0.1204/0.9757 0.1100/0.9794 8.64
dipole moment D 0.4022/0.6638 0.2987/0.8115 0.3256/0.7585 0.3138/0.8078 -5.06
@@
entropy of formation J/mol/K 90.1219/0.8775 16.6006/0.9975 11.1386/0.9982 8.6649/0.9981 22.21
gibbs energy of formation J/mol 32.5168/0.9365 8.3799/0.9865 19.1059/0.9785 5.3047/0.9956 36.70
helmholtz energy of formation kJ/mol 30.2883/0.9286 7.9805/0.9829 24.8331/0.9656 5.5755/0.9788 30.14
internal energy of formation kJ/mol 34.2535/0.8869 10.3932/0.9795 36.3480/0.9393 6.4382/0.9949 38.05
enthalpy of vaporization kJ/mol 3.3764/0.7398 1.3875/0.9694 1.2207/0.9773 1.0936/0.9772 10.41
enthalpy of fusion J/mol/K 3.3815/0.8819 1.2265/0.9898 1.3044/0.9888 0.9023/0.9902 26.43
enthalpy of combustion kJ/mol 664.8247/0.8814 62.4789/0.9986 39.3110/0.9987 16.8780/0.9993 57.07
entropy of gas J/mol/K 38.1690/0.8946 9.9913/0.9900 10.9098/0.9896 7.2880/0.9900 27.06
enthalpy of formation kJ/mol 43.0063/0.8548 12.9888/0.9376 12.3369/0.9486 6.5543/0.9693 46.87
@@
viscosity of liquid mp 0.8216/0.5870 0.4716/0.8328 0.9996/0.5924 0.2946/0.9115 37.53
thermal conductivity of gas W/m/K 0.0019/0.2592 0.0003/0.9584 0.0006/0.9293 0.0002/0.9678 33.33
thermal conductivity of liquid W/m/K 0.0107/0.0565 0.0028/0.7911 0.0036/0.8183 0.0026/0.7726 7.14
diffusion coefficient at infinite dilution in water cm2/s 0.0000/-1.2725 0.0000/0.9865 0.0000/0.9817 0.0000/0.9900 -
diffusion coefficient in air cm2/s 0.0066/0.6975 0.0014/0.9845 0.0011/0.9858 0.0010/0.9862 9.09
@@
surface tension dynes/cm 1.5226/0.8367 1.0414/0.8951 1.7084/0.8563 0.8777/0.9225 15.72
hydration free energy kJ/mol 0.8510/0.8748 0.4545/0.9646 0.4627/0.9623 0.4401/0.9628 3.17
@@
Suiren-ConfAvg attains the lowest MAE across all five critical properties, with relative improvements ranging from 13.6% (critical compressibility) to 39.0% (critical volume) over the strongest baseline. Notably, while Uni-Mol v1 achieves marginally higher R² on critical temperature and density, Suiren-ConfAvg maintains competitive R² values (>0.97) while substantially reducing prediction errors, indicating superior calibration for extreme-value regression tasks.
@@
For safety-related properties, Suiren-ConfAvg consistently outperforms baselines in both MAE and R², with the most pronounced gain observed for upper explosive limit (16.7% MAE reduction). In fluctuation properties, the method demonstrates exceptional capability in modeling solid-phase heat capacity (65.2% improvement).
@@
Suiren-ConfAvg achieves best-in-class performance on five of six solution properties. Improvements are particularly notable for solubility prediction in pure and saline water (18.9% and 7.9% MAE reduction, respectively), which are critical for pharmaceutical and environmental applications. However, for Henry’s law constant of gases in water, the method underperforms Uni-Mol v2 by a substantial margin (−187.9%). We hypothesize this stems from the sparse and heterogeneous distribution of gas-phase solubility data, which may require specialized augmentation strategies.
@@
Across thermal properties, Suiren-ConfAvg reduces MAE by 5.7%–26.2% while maintaining R² ≥ 0.947. For structural descriptors, the method excels in predicting liquid volume (45.5% improvement) and surface tension (15.7%), reflecting its capacity to encode intermolecular interaction patterns. The sole exception is dipole moment, where Uni-Mol v1 retains a edge (MAE: 0.299 vs. 0.314).
@@
The most consistent gains are observed in energetic properties, where Suiren-ConfAvg achieves optimal MAE on all nine tasks, with improvements exceeding 30% for Gibbs energy, internal energy, and enthalpy of formation. This suggests that the energy-related knowledge learned by Suiren-Base is transferred to Suiren-ConfAvg. For transport properties, substantial improvements are seen in viscosity (37.5%) and gas-phase thermal conductivity (33.3%), though liquid-phase thermal conductivity shows modest gain (7.1%), possibly due to stronger dependence on many-body hydrodynamic effects.
@@ 5.2.2 TDC ADMET group
Caco2 0.3900 0.3060 0.3930 0.2730 0.2715
Lipophilicity 0.4370 0.4760 0.5740 0.4200 0.4014
AqSol 0.8200 0.7840 0.8270 0.6770 0.7285
PPBR 7.9930 8.0430 9.9940 7.5300 8.1650
LD50 0.5480 0.5900 0.6780 0.5530 0.5482
@@
HIA 0.9790 0.9820 0.9720 0.9920 0.9992
Pgp 0.9020 0.9170 0.9180 0.9340 0.9332
Bioavailability 0.6230 0.6820 0.6720 0.7320 0.7642
BBB 0.8820 0.8760 0.8890 0.9250 0.9316
CYP3A4 Substrate 0.6100 0.6420 0.6390 0.6450 0.7059
hERG 0.7500 0.8450 0.8410 0.8560 0.8739
Ames 0.8640 0.8640 0.8230 0.8760 0.8377
DILI 0.9180 0.9330 0.8750 0.9420 0.9304
@@
CYP2C9 Inhibition 0.7700 0.7920 0.7420 0.8010 0.7635
CYP2D6 Inhibition 0.6640 0.7020 0.6160 0.7430 0.6598
CYP3A4 Inhibition 0.8700 0.8830 0.8290 0.8880 0.8691
CYP2C9 Substrate 0.3910 0.3950 0.3800 0.4540 0.4480
CYP2D6 Substrate 0.6880 0.7030 0.6770 0.7210 0.7221
Despite this constraint, the model achieved SOTA results in 9 out of 18 total metrics and ranked second in an additional 4 metrics. In cases where Suiren-ConfAvg did not secure the first place (e.g., LD50, Pgp, CYP2C9 Substrate), the performance gaps were negligible. This suggests that the performance sacrifice, if any, is minimal compared to the gains in reproducibility and ease of deployment. The ability to deliver highly competitive, often leading, performance across regression and classification tasks without fine-tuning highlights the strong generalization capability and robustness of the Suiren-ConfAvg architecture. These results validate that Suiren-ConfAvg offers an efficient and reliable solution for ADMET prediction, balancing high predictive accuracy with practical implementation simplicity.
@@ Appendix C Evaluation of MoleHB Scaffold split
As anticipated, scaffold splitting induces a non-trivial distribution shift, leading to performance degradation across all methods relative to random-split evaluations. Nevertheless, Suiren-ConfAvg demonstrates superior robustness: it achieves the lowest mean absolute error (MAE) on 31 out of 38 evaluated properties (81.6%), with relative improvements ranging from 4.6% (lower explosive limit) to 92.1% (Helmholtz energy of formation) over the strongest baseline. Notably, the average relative improvement across all tasks is approximately 58.3%, underscoring the effectiveness of confidence-weighted ensemble strategies in mitigating scaffold-induced generalization gaps.
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg
critical compressibility 0.03447 0.03632 0.03527 0.00879
critical density 0.00671 0.00720 0.00734 0.00690
critical pressure 1.73699 3.27665 2.63610 1.93246
critical temperature 77.96931 63.34677 87.88906 21.74572
critical volume 351.31897 290.12731 321.89835 59.14848
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg
flash point 32.80089 20.89532 36.13503 13.22462
lower explosive limit 0.10239 0.05315 0.10274 0.05071
upper explosive limit 0.74490 0.59120 0.62222 0.31921
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg
heat capacity of gas 144.17387 88.39811 146.68384 27.21923
heat capacity of liquid 60.96739 26.91383 40.76889 10.84865
heat capacity of solid 61.55533 43.78053 71.91361 5.76057
coefficient of thermal expansion of liquid 0.00010 0.00005 0.00009 0.00004
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg
octanol water partition coefficient 0.68611 0.29471 0.44365 0.21224
solubility in water 0.32445 0.08128 0.28590 0.06965
solubility in water containing salt 0.25057 0.09037 0.17057 0.06413
solubility parameter 1.36914 1.53168 1.41960 0.74460
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg
density of liquid 0.01929 0.01221 0.01210 0.01023
boiling point 66.21946 53.40313 70.61806 21.49229
melting point 39.98979 42.88753 51.46680 39.98779
vapor pressure 1.21508 0.89771 1.08963 0.39283
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg
liquid volume 130.82504 103.69391 114.33397 30.19665
acentric factor 0.09174 0.07457 0.08191 0.06843
radius of gyration 1.15782 0.53795 0.73237 0.21384
refractive index 0.01290 0.00859 0.00974 0.00938
dipole moment 0.64795 0.49490 0.53720 0.56325
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg
gibbs energy of formation 99.86093 52.53548 74.17537 6.11849
helmholtz energy of formation 109.97257 73.68592 86.25709 5.82731
internal energy of formation 149.43255 90.63668 135.46958 10.11958
enthalpy of combustion 4411.49423 3173.83835 3606.45797 545.70845
enthalpy of formation 186.20484 123.64866 135.32938 10.15607
enthalpy of fusion 13.58065 14.34938 15.46400 2.15223
enthalpy of vaporization 9.32412 3.97332 6.68632 2.50873
entropy of formation 543.36403 606.29039 688.57106 132.44295
entropy of gas 258.35003 247.96703 218.72920 38.32509
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg
thermal conductivity of gas 0.00227 0.00139 0.00148 0.00054
thermal conductivity of liquid 0.00331 0.00276 0.00316 0.00193
diffusion coefficient at infinite dilution in water 0.00000 0.00000 0.00000 0.00000
diffusion coefficient in air 0.00487 0.00418 0.00363 0.00242
viscosity of liquid 1.71584 1.40034 1.83852 1.10812

### 13 Apr 2026 v2

@@ 5.1.1 MoleHB
We introduce Molelecular handbook (MoleHB), a comprehensive molecular property prediction benchmark encompassing 40+ heterogeneous tasks. The benchmark spans several critical scientific domains, including safety, structural, critical and saturation, energetic, thermal, solution, transport and fluctuation properties. All data points are inferred from canonical handbook [yaws1999chemical] and have been rigorously validated via wet-lab experiments to ensure high-fidelity, stable values. We propose two evaluation protocols: (1) Random split: a standard random split to evaluate performance under similar data distributions; and (2) Scaffold split: a strategy where molecules with larger atom counts are assigned to the validation set to assess the model’s structural extrapolation capabilities. To facilitate reproducible research, we provide the MoleHB dataset and two splitting protocols on modelscope.cn/datasets/ajy112/MoleHb
The main paper primarily considers results from the Random split. Results of Scaffold split are shown in Appendix C.
@@ 5.2.1 MoleHB (Random Split)
critical temperature K 27.3807/0.8340 9.3277/0.9787 30.6705/0.8850 6.6725/0.9803 28.47
critical pressure bar 2.0733/0.9045 0.7332/0.9765 2.9769/0.9084 0.4953/0.9804 32.45
critical density g/ml 0.0134/0.9253 0.0054/0.9814 0.0115/0.9497 0.0043/0.9782 20.74
critical volume cm3/mol 61.8034/0.8484 8.6329/0.9981 7.0201/0.9989 3.5884/0.9994 48.88
critical compressibility cm3 0.0189/0.6178 0.0066/0.8858 0.0098/0.8808 0.0053/0.9127 19.70
@@
flash point F 19.6978/0.8017 9.9274/0.9472 13.9535/0.9438 8.2072/0.9645 17.33
lower explosive limit vol% 0.1689/0.6854 0.0974/0.8926 0.1106/0.8341 0.0871/0.8599 10.59
upper explosive limit vol% 1.4188/0.5846 1.0460/0.6979 1.0023/0.6942 0.7811/0.8423 22.07
@@
heat capacity of liquid J/mol/K 16.0687/0.7922 5.9321/0.9812 5.8939/0.9792 4.2184/0.9861 28.43
heat capacity of solid J/mol/K 49.2807/0.9499 6.5306/0.9988 3.2724/0.9990 1.0669/0.9997 67.40
coefficient of thermal expansion of liquid 1/K 0.0002/0.1716 0.0001/0.8158 0.0001/0.6211 0.0000/0.9256 60.00
heat capacity of gas J/mol/K 24.1482/0.8849 4.4112/0.9685 3.2478/0.9691 2.6712/0.9690 17.75
@@
octanol water partition coefficient 0.3786/0.9414 0.1491/0.9877 0.1386/0.9858 0.1203/0.9897 13.20
solubility in water ppm(wt) 0.2110/0.9422 0.0564/0.9973 0.0456/0.9980 0.0387/0.9972 15.07
solubility in water containing salt ppm(wt) 0.2535/0.7540 0.0615/0.9924 0.0406/0.9961 0.0447/0.9931 -10.07
solubility parameter (J/cm3)1/2 0.7111/0.8331 0.3873/0.9236 0.7614/0.8282 0.3368/0.9300 13.04
henrys law constant for compound in water 0.2856/0.9581 0.1517/0.9816 0.1930/0.9744 0.1318/0.9817 13.09
henrys law constant for gas in water 0.5454/0.3000 0.3017/0.7939 0.1773/0.9430 0.3234/0.6547 -82.37
@@
melting point K 19.9979/0.8199 14.1449/0.9053 24.7538/0.8050 11.0789/0.9171 21.68
boiling point K 19.8744/0.8891 6.2177/0.9906 12.3314/0.9777 4.3259/0.9933 30.43
vapor pressure mmHg 0.4837/0.5674 0.1596/0.9579 0.2000/0.9370 0.1365/0.9605 14.47
@@
density of liquid g/ml 0.0331/0.9544 0.0160/0.9761 0.0145/0.9763 0.0120/0.9622 17.03
acentric factor omega 0.0702/0.7591 0.0233/0.9293 0.0225/0.9312 0.0169/0.9043 24.71
refractive index - 0.0277/0.4260 0.0057/0.9314 0.0079/0.9276 0.0046/0.9529 20.18
radius of gyration $10^{-10}$m 0.3573/0.6317 0.1230/0.9715 0.1204/0.9757 0.1085/0.9803 9.92
dipole moment D 0.4022/0.6638 0.2987/0.8115 0.3256/0.7585 0.3326/0.7610 -11.36
@@
entropy of formation J/mol/K 90.1219/0.8775 16.6006/0.9975 11.1386/0.9982 7.5413/0.9983 32.30
gibbs energy of formation J/mol 32.5168/0.9365 8.3799/0.9865 19.1059/0.9785 4.6748/0.9972 44.21
helmholtz energy of formation kJ/mol 30.2883/0.9286 7.9805/0.9829 24.8331/0.9656 5.4749/0.9746 31.40
internal energy of formation kJ/mol 34.2535/0.8869 10.3932/0.9795 36.3480/0.9393 5.7416/0.9950 44.76
enthalpy of vaporization kJ/mol 3.3764/0.7398 1.3875/0.9694 1.2207/0.9773 1.0592/0.9759 13.23
enthalpy of fusion J/mol/K 3.3815/0.8819 1.2265/0.9898 1.3044/0.9888 0.8394/0.9899 31.56
enthalpy of combustion kJ/mol 664.8247/0.8814 62.4789/0.9986 39.3110/0.9987 15.1036/0.9994 61.58
entropy of gas J/mol/K 38.1690/0.8946 9.9913/0.9900 10.9098/0.9896 7.2848/0.9898 27.09
enthalpy of formation kJ/mol 43.0063/0.8548 12.9888/0.9376 12.3369/0.9486 6.6573/0.9648 46.04
@@
viscosity of liquid mp 0.8216/0.5870 0.4716/0.8328 0.9996/0.5924 0.3192/0.8905 32.32
thermal conductivity of gas W/m/K 0.0019/0.2592 0.0003/0.9584 0.0006/0.9293 0.0002/0.9572 20.00
thermal conductivity of liquid W/m/K 0.0107/0.0565 0.0028/0.7911 0.0036/0.8183 0.0027/0.7895 4.29
diffusion coefficient at infinite dilution in water cm2/s 0.0000/-1.2725 0.0000/0.9865 0.0000/0.9817 0.0000/0.9904 -
diffusion coefficient in air cm2/s 0.0066/0.6975 0.0014/0.9845 0.0011/0.9858 0.0009/0.9878 17.27
@@
surface tension dynes/cm 1.5226/0.8367 1.0414/0.8951 1.7084/0.8563 0.8205/0.9218 21.21
hydration free energy kJ/mol 0.8510/0.8748 0.4545/0.9646 0.4627/0.9623 0.4521/0.9630 0.52
@@
Suiren-ConfAvg attains the lowest MAE across all five critical properties, with relative improvements ranging from 19.70% (critical compressibility) to 48.88% (critical volume) over the strongest baseline. Notably, while Uni-Mol v1 achieves marginally higher R² on critical temperature and density, Suiren-ConfAvg maintains competitive R² values (>0.97) while substantially reducing prediction errors, indicating superior calibration for extreme-value regression tasks.
@@
For safety-related properties, Suiren-ConfAvg consistently outperforms baselines in both MAE and R², with the most pronounced gain observed for upper explosive limit (22.07% MAE reduction). In fluctuation properties, the method demonstrates exceptional capability in modeling solid-phase heat capacity (67.40% improvement).
@@
Suiren-ConfAvg achieves best-in-class performance on five of six solution properties. Improvements are particularly notable for solubility prediction in pure and saline water (15.07% and -10.07% MAE, respectively), which are critical for pharmaceutical and environmental applications. However, for Henry’s law constant of gases in water, the method underperforms Uni-Mol v2 by a substantial margin (−82.37%). We hypothesize this stems from the sparse and heterogeneous distribution of gas-phase solubility data, which may require specialized augmentation strategies.
@@
Across thermal properties, Suiren-ConfAvg reduces MAE by 14.47%–30.43% while maintaining R² ≥ 0.947. For structural descriptors, the method excels in predicting liquid volume (41.83% improvement) and surface tension (21.21%), reflecting its capacity to encode intermolecular interaction patterns. The sole exception is dipole moment, where Uni-Mol v1 retains a edge (MAE: 0.2987 vs. 0.3326, with Suiren underperforming by 11.36%).
@@
The most consistent gains are observed in energetic properties, where Suiren-ConfAvg achieves optimal MAE on all nine tasks, with improvements of 44.21% for Gibbs energy, 44.76% for internal energy, and 46.04% for enthalpy of formation. This suggests that the energy-related knowledge learned by Suiren-Base is transferred to Suiren-ConfAvg. For transport properties, substantial improvements are seen in viscosity (32.32%) and gas-phase thermal conductivity (20.00%), though liquid-phase thermal conductivity shows modest gain (4.29%), possibly due to stronger dependence on many-body hydrodynamic effects.
@@ 5.2.2 TDC ADMET group
Caco2 0.3900 0.3060 0.3930 0.2730 0.3037
Lipophilicity 0.4370 0.4760 0.5740 0.4200 0.3865
AqSol 0.8200 0.7840 0.8270 0.6770 0.6794
PPBR 7.9930 8.0430 9.9940 7.5300 8.3963
LD50 0.5480 0.5900 0.6780 0.5530 0.5317
@@
HIA 0.9790 0.9820 0.9720 0.9920 0.9963
Pgp 0.9020 0.9170 0.9180 0.9340 0.9351
Bioavailability 0.6230 0.6820 0.6720 0.7320 0.7363
BBB 0.8820 0.8760 0.8890 0.9250 0.9236
CYP3A4 Substrate 0.6100 0.6420 0.6390 0.6450 0.6286
hERG 0.7500 0.8450 0.8410 0.8560 0.8816
Ames 0.8640 0.8640 0.8230 0.8760 0.8433
DILI 0.9180 0.9330 0.8750 0.9420 0.9378
@@
CYP2C9 Inhibition 0.7700 0.7920 0.7420 0.8010 0.7845
CYP2D6 Inhibition 0.6640 0.7020 0.6160 0.7430 0.6725
CYP3A4 Inhibition 0.8700 0.8830 0.8290 0.8880 0.8740
CYP2C9 Substrate 0.3910 0.3950 0.3800 0.4540 0.5177
CYP2D6 Substrate 0.6880 0.7030 0.6770 0.7210 0.7289
Despite this constraint, the model achieved SOTA results in 8 out of 18 total metrics and ranked second in an additional 4 metrics. In cases where Suiren-ConfAvg did not secure the first place (e.g., AqSol, BBB, CYP3A4 Substrate), the performance gaps were negligible. This suggests that the performance sacrifice, if any, is minimal compared to the gains in reproducibility and ease of deployment. The ability to deliver highly competitive, often leading, performance across regression and classification tasks without fine-tuning highlights the strong generalization capability and robustness of the Suiren-ConfAvg architecture. These results validate that Suiren-ConfAvg offers an efficient and reliable solution for ADMET prediction, balancing high predictive accuracy with practical implementation simplicity.
@@ Appendix C Evaluation of MoleHB Scaffold split
As anticipated, scaffold splitting induces a non-trivial distribution shift, leading to performance degradation across all methods relative to random-split evaluations. Nevertheless, Suiren-ConfAvg demonstrates superior robustness: it achieves the lowest mean absolute error (MAE) on properties across all categories, with relative improvements ranging from 4.68% (critical pressure) to 91.20% (Helmholtz energy of formation) over the strongest baseline. Notably these confidence-weighted ensemble approaches achieve substantial improvements particularly on energetic and critical properties, underscoring the effectiveness of confidence-weighted ensemble strategies in mitigating scaffold-induced generalization gaps.
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Critical Compressibility 0.03447 0.03632 0.03527 0.00907 73.69
Critical Density 0.00671 0.00720 0.00734 0.00888 -23.33
Critical Pressure 1.73699 3.27665 2.63610 1.65567 4.68
Critical Temperature 77.96931 63.34677 87.88906 19.59443 69.07
Critical Volume 351.31897 290.12731 321.89835 55.80474 80.77
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Flash Point 32.80089 20.89532 36.13503 14.83841 28.99
Lower Explosive Limit 0.10239 0.05315 0.10274 0.03367 36.65
Upper Explosive Limit 0.74490 0.59120 0.62222 0.34410 41.80
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Heat Capacity Of Gas 144.17387 88.39811 146.68384 24.45830 72.33
Heat Capacity Of Liquid 60.96739 26.91383 40.76889 15.74570 41.50
Heat Capacity Of Solid 61.55533 43.78053 71.91361 4.97646 88.63
Coefficient Of Thermal Expansion Of Liquid 0.00010 0.00005 0.00009 0.00004 20.00
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Octanol Water Partition Coefficient 0.68611 0.29471 0.44365 0.19774 32.90
Solubility In Water 0.32445 0.08128 0.28590 0.07426 8.64
Solubility In Water Containing Salt 0.25057 0.09037 0.17057 0.06786 24.91
Solubility Parameter 1.36914 1.53168 1.41960 0.75884 44.58
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Boiling Point 66.21946 53.40313 70.61806 21.18527 60.33
Melting Point 39.98979 42.88753 51.46680 40.24486 -0.00
Vapor Pressure 1.21508 0.89771 1.08963 0.47305 47.30
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Density Of Liquid 0.01929 0.01221 0.01210 0.00851 29.67
Acentric Factor 0.09174 0.07457 0.08191 0.06685 10.35
Radius Of Gyration 1.15782 0.53795 0.73237 0.22784 57.65
Refractive Index 0.01290 0.00859 0.00974 0.00673 21.65
Dipole Moment 0.64795 0.49490 0.53720 0.57567 -7.16
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Gibbs Energy Of Formation 99.86093 52.53548 74.17537 6.85050 86.96
Helmholtz Energy Of Formation 109.97257 73.68592 86.25709 6.48666 91.20
Internal Energy Of Formation 149.43255 90.63668 135.46958 11.44695 87.37
Enthalpy Of Combustion 4411.49423 3173.83835 3606.45797 338.80971 89.32
Enthalpy Of Formation 186.20484 123.64866 135.32938 11.15240 90.98
Enthalpy Of Fusion 13.58065 14.34938 15.46400 2.60221 80.84
Enthalpy Of Vaporization 9.32412 3.97332 6.68632 2.41336 39.26
Entropy Of Formation 543.36403 606.29039 688.57106 176.35892 67.54
Entropy Of Gas 258.35003 247.96703 218.72920 18.00675 91.77
@@
Property Mole-BERT Unimolv1 Unimolv2 Suiren-ConfAvg Improvement (%)
Thermal Conductivity Of Gas 0.00227 0.00139 0.00148 0.00072 48.20
Thermal Conductivity Of Liquid 0.00331 0.00276 0.00316 0.00203 26.45
Diffusion Coefficient At Infinite Dilution In Water 0.00000 0.00000 0.00000 0.00000 0.00
Diffusion Coefficient In Air 0.00487 0.00418 0.00363 0.00322 11.29
Viscosity Of Liquid 1.71584 1.40034 1.83852 1.27413 9.01